### v1

# Efficient Inference for Large Reasoning Models: A Survey

Abstract
Large Reasoning Models (LRMs) significantly improve the reasoning ability of Large Language Models (LLMs) by learning to reason, exhibiting promising performance in complex task-solving. However, their deliberative reasoning process leads to inefficiencies in token usage, memory consumption, and inference time. Thus, this survey provides a review of efficient inference methods designed specifically for LRMs, focusing on mitigating token inefficiency while preserving the reasoning quality. The overview structure of this paper is shown in Figure 2. First, we introduce a taxonomy to group the recent methods into two main categories: (a) explicit compact Chain-of-Thought (CoT), which reduces tokens while keeping the explicit reasoning structure, and (b) implicit latent CoT, which encodes reasoning steps within hidden representations instead of explicit tokens. Meanwhile, we discuss their strengths and weaknesses. Then, we conduct empirical analyses on existing methods from performance and efficiency aspects. Besides, we present open challenges in this field, including human-centric controllable reasoning, trade-off between interpretability and efficiency of reasoning, ensuring safety of efficient reasoning, and broader applications of efficient reasoning. In addition, we highlight key insights for enhancing LRMs’ inference efficiency via techniques such as model merging, new architectures, and agent routers. We hope this work serves as a valuable guide, helping researchers overcome challenges in this vibrant field [^1] https://github.com/yueliu1999/Awesome-Efficient-Inference-for-LRMs.
1 Introduction
Large Language Models (LLMs), which are trained to provide quick and intuitive responses, have exhibited great success in various fast-thinking applications like ChatBot (OpenAI, 2022). Differently, slow-thinking scenarios like math problem-solving (Olympiad, 2025) or research (OpenAI, 2025a) require the models to conduct analytical and deliberative reasoning before providing final responses. To tackle these challenges, Large Reasoning Models (LRMs) such as OpenAI o1/o3 (Jaech et al., 2024; OpenAI, 2025) and DeepSeek R1 (Guo et al., 2025) are developed by guiding the model to learn to reason.
Although effective, the intermediate reasoning process of LRMs is highly resource-intensive, learning to three challenges: (1) significant token consumption, (2) high memory overhead, and (3) increased inference time. These problems not only increase the inference cost of the service companies but also degrade the experience of the users. Therefore, efficient inference for LRMs has become an urgent and crucial direction. Since thinking tokens are treated like regular output tokens without cost differentiation, previous efforts in inference efficiency of regular LLMs, e.g., model compression (Wang et al., 2024b), efficient model design (Nie et al., 2025), and system-level optimization (Liu et al., 2024a), can alleviate problems (2) and (3). These methods are comprehensively studied (Zhou et al., 2024) and not specially designed for LRMs. Therefore, this paper focuses on the challenge (1): token inefficiency.
Figure 1: Overview Structure of this Survey.
Figure 2: Overview of Taxonomy.
To this end, we conduct a comprehensive survey of recent efficient inference methods designed specifically for LRMs, aiming at improving thinking token efficiency while preserving reasoning quality. Specifically, we present a hierarchical taxonomy that first categorizes recent approaches into two classes. As shown in Figure 2, it contains (a) the explicit compact CoT, which reduces the number of thinking tokens while maintaining explicit reasoning structure, and (b) the implicit latent CoT, which encodes reasoning steps within hidden representations instead of explicit tokens. In addition, for the explicit compact CoT, we further summarize three sub-categories: (a.1) CoT compression, (a.2) CoT preference optimization, and (a.3) reward-based CoT conciseness. Then, we analyze the characteristics of these categories and discuss their strengths and weaknesses from the aspects of reasoning quality and efficiency. Besides, we conduct a comprehensive empirical study on the existing methods from the perspectives of performance and efficiency. Besides, we identify four open challenges regarding the inference efficiency of LRMs, including human-centric controllable reasoning, the trade-off between efficiency and interoperability of reasoning, ensuring the safety of efficient reasoning, and broader applications of efficient LRMs beyond math and code. Last, we highlight several potential technical solutions for further improvement of current methods, i.e., model merging, new architectures, and agent routers. We hope that this survey helps researchers and engineers further improve efficient inference for LRMs. The main contributions of this paper are summarized as follows.
  1. We conduct a comprehensive paper review of current methods of efficient inference for LRMs with a hierarchical taxonomy and strength & weakness discussion.
  2. We empirically study recent methods from performance and efficiency views and summarize 4 challenges from user control, interpretability, safety, and application aspects.
  3. We highlight technical insights in further improvement of existing methods from the perspectives of model merging, non-autoregressive architectures, and agent routers
2 Background
This section first introduces the background of large reasoning models and then highlights the efficiency challenges in the inference phase of large reasoning models.
2.1 Large Reasoning Model
Large Reasoning Models (LRMs) extend the capabilities of Large Language Models (LLMs) by incorporating explicit intermediate tokens that represent reasoning processes, enabling more structured logical reasoning and effective complex problem-solving. LRMs mimic the way humans approach complex problems by first thinking before providing an answer. When faced with a difficult question, they do not immediately respond with an answer; instead, they analyze the problem, break it down into smaller steps, explore different solution paths, and verify their reasoning before arriving at a conclusion. The o1 series (Jaech et al., 2024) from OpenAI, released in late 2024, marked a significant breakthrough in AI reasoning capabilities, which integrates reinforcement learning and “Chain of-Thought” prompting (Wei et al., 2022) techniques. Following this, OpenAI released o3 (OpenAI, 2025), an upgraded version of o1, allowing it to achieve PhD-level performance in mathematics, science, and programming. Notable DeepSeek’s R1 (Guo et al., 2025) stands out for being open-sourced, with transparent thinking process tokens, which sets it apart from other proprietary LRMs like o1/o3, where the internal reasoning steps are less accessible. However, since LRMs need to generate numerous intermediate thinking tokens before arriving at final answers, they are significantly less efficient and more expensive compared to regular LLMs. This added complexity in processing demands more computational resources and time.
Types Methods Training Strategy Model Application
Explicit Compact CoT SoT (Aytes et al., 2025) Prompt Qwen-2.5-7B/14B/32B Math, Commonsense, Logic, Scientific, Medical
Constrained-CoT (Nayab et al., 2024) Prompt LLaMA-2-70B, Falcon-40B Math
CoD (Xu et al., 2025b) Prompt GPT-4o, Claude 3.5 Sonnet Math, Commonsense, Symbolic Reasoning
TALE-EP (Han et al., 2024) Prompt LLaMA-3.1-8B-Instruct Math
Meta-Reasoner (Sui et al., 2025) Prompt GPT-4o, GPT-4o-mini, Gemini-Exp-1206 Math, Scientific
SOLAR (Li et al., 2025) SFT Qwen2VL-7B-Instruct Math
C3oT (Kang et al., 2024) SFT LLaMA-2-Chat -7B & -13B Math, Commonsense
TokenSkip (Xia et al., 2025) SFT LLaMA-3.1-8B-Instruct, Qwen2.5- 14B-Instruct Math
InftyThink (Yan et al., 2025) SFT Qwen2.5-14B/32B, Qwen2.5-Math-1.5B/7B, LLaMA-3.1-8B Math, Scientific
LightThinker (Zhang et al., 2025) SFT DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-LLaMA-8B Language Understanding, Math, Scientific, Commonsense, Logic
CoT-Valve (Ma et al., 2025) SFT QwQ-32B-Preview, DeepSeek-R1-Distill-LLaMA-8B, LLaMA-3.1-8B, LLaMA-3.2-1B, Qwen32B-Instruct Math
Distill System 2 (Yu et al., 2024) SFT LLaMA-2-70B-chat Math, Commonsense, Coin Flip
SF (Munkhbat et al., 2025) SFT LLaMA-3.2-3B, Gemma2-2B , Qwen2.5-3B , Qwen2.5-Math-1.5B, DeepSeekMath-7B Math
Skip Steps (Liu et al., 2024c) SFT LLaMA2-7b, Phi-3-mini Math, Logic
VARR (Jang et al., 2024) SFT Mistral 7B, Llama3.2 1B/3B Math, Commonsense
DAST (Shen et al., 2025b) SimPO DS-R1-Distill-Qwen-7B, DS-R1-Distill-Qwen-32B Math
TALE-PT (Han et al., 2024) SFT, DPO LLaMA-3.1-8B-Instruct Math
Kimi k1.5 (Kimi Team et al., 2025) RL Kimi k1.5 Multimodal Understanding, Math, Code
O1-Pruner (Luo et al., 2025) RL Marco-o1-tB, QwQ-32B Math
MRT (Qu et al., 2025) RL DeepSeek-R1-Distill-Qwen-32B Math
(Arora & Zanette, 2025) RL DS-R1-Distill-Qwen-1.5B, DS-R1-Distill-Qwen-7B Math
Claude 3.7 (Anthropic, 2025) RL Unknown Math, Code, Agent
L1 (Aggarwal & Welleck, 2025) RL Qwen-Distilled-R1-1.5B Language Understanding, Logic, Math
SPIRIT (Cui et al., 2025) RL LLaMA3-8B-Instruct, Qwen2.5- 7B-Instruct Math
IBPO (Yu et al., 2025) RL LLaMA-3.1-8B Math
Implicit Latent CoT ICoT-KD (Deng et al., 2023) SFT GPT-2 Small/Medium Math
CODI (Shen et al., 2025c) SFT GPT-2 Small, LLaMA-3.2-1B Math
ICoT-SI (Deng et al., 2024) SFT GPT-2 Small/Medium, Phi-3 3.8B, Mistral 7B Math
COCONUT (Hao et al., 2024) SFT GPT-2 Math
CCoT (Cheng & Van Durme, 2024) SFT LLaMA2-7B-Chat Math, Logic
Heima (Shen et al., 2025a) SFT LLaVA-CoT, LLaMA-3.1-8B-Instruct Multimodal Reasoning
Token assorted (Su et al., 2025) SFT LLaMA-3.2-1B, LLaMA-3.2-3B, LLaMA-3.1-8B Agentic Planning, Logic, Math.
SoftCoT (Xu et al., 2025c) SFT LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct Math, Commonsense, Symbolic Reasoning
Table 1: A taxonomy of efficient inference methods for Large Reasoning Models.
2.2 Efficiency Challenge in LRM Inference
A key driver of LRMs’ remarkable reasoning capabilities is the scaling of inference-time compute, which enables complex reasoning through long CoTs (Chen et al., 2025; Guo et al., 2025; Jaech et al., 2024; Muennighoff et al., 2025; Liu et al., 2025a). Compared to standard short CoTs (Wei et al., 2022), which are often shallow, heuristic-driven, and less generalizable (Sprague et al., 2025), long CoTs empower LRMs to tackle complex tasks such as advanced mathematics (Xu et al., 2025a) and medical question answering (Huang et al., 2025). However, this shift has also introduced the phenomenon of overthinking, where LRMs consume excessive inference tokens and reasoning steps even for simple problems, yielding only marginal performance improvements (Ma et al., 2024; Chen et al., 2024; Wu et al., 2025c). In real-world applications such as software engineering agents, overthinking has been found to negatively correlate with issue resolution rates (Cuadron et al., 2025). Moreover, LRMs’ reliance on inference-time scaling exposes them to overthinking attacks, where adversarial actors inject benign yet computationally intensive decoy problems (e.g., Sudoku puzzles) into the context for retrieval-augmented question answering, triggering substantial computational overhead (Kumar et al., 2025).
Toward practical real-world deployment, optimizing the token efficiency of LRMs without compromising effectiveness remains an underexplored challenge. This paper presents a systematic investigation into recent advances in token-efficient LRMs, examining their underlying approaches, empirical effectiveness, and implications for future research.
3 Landscape of LRM Efficient Inference Research
This section surveys the current landscape of research on token-efficient LRM inference, which can be broadly categorized into two approaches: (1) explicit compact CoT, where explicit instructions, rewards, or budget constraints are introduced to encourage shorter reasoning chains over long CoTs (Section 3.1); and (2) implicit latent CoT, which compresses explicit long CoTs into compact, continuous reasoning states (Section 3.2). The taxonomy of recent efficient inference methods is shown in Table 1.
3.1 Explicit Compact CoT
Recent research has focused on developing methods to create more compact reasoning paths while preserving accuracy through various techniques, including (1) CoT compression, (2) fine-tuning for compact reasoning, and (3) reward-based incentivization.
CoT Compression. Succinct CoT representations streamline inference while preserving solution quality. Constrained-CoT (Nayab et al., 2024) and CoD (Xu et al., 2025b) confine intermediate reasoning to essential steps, ensuring brevity without losing critical information. Beyond simple compression, Sketch-of-Thought (SoT) (Aytes et al., 2025) uses a smaller “router” model to prompt the main LLM to generate sketches of reasoning, offering a concise yet cognitively inspired overview. InftyThink (Yan et al., 2025) decomposes complex tasks into bounded-length segments, creating intermediate summaries at each step. Other frameworks adapt their compression mechanisms in real time. LightThinker (Zhang et al., 2025) introduces special tokens that trigger the model to dynamically compress its ongoing thought process, reducing redundancy. TALE-EP (Han et al., 2024) dynamically adjusts the allotted reasoning tokens depending on task complexity, while Meta-Reasoner (Sui et al., 2025) applies a contextual multi-armed bandit to optimize efficiency.
Fine-Tuning on Compact Reasoning Chains. Fine-tuning on compact reasoning data enables LRMs to internalize efficient inference behaviors while keeping performance. C3oT (Kang et al., 2024) leverages an LLM to generate condensed versions of long CoTs, preserving essential structure before jointly training models on both full and compressed chains. Skip Steps (Liu et al., 2024c) curates expert-validated answers with condensed steps and fine-tunes LLMs to mimic these concise reasoning paths. SOLAR (Li et al., 2025) fine-tunes LLMs using datasets annotated for both correctness and the effectiveness of the underlying task-specific reasoning topology, encouraging minimal yet complete logic flows. VARR (Jang et al., 2024) identifies redundant sentences in CoTs by analyzing their contribution to answer correctness, then fine-tunes models on the distilled, non-redundant reasoning processes. TokenSkip (Xia et al., 2025) prunes reasoning chains token-by-token based on importance, followed by fine-tuning across various compression ratios to balance brevity and precision. From a parameter space perspective, TALE-EP (Han et al., 2024) enhances token-budget awareness via SFT and direct preference optimization (DPO). CoT-Valve (Ma et al., 2025) discovers a latent direction that controls reasoning length, enabling models to flexibly adjust their level of detail based on task demands.
Reward-Based Incentivization. A growing body of work introduces explicit reward signals to reduce unnecessary CoT complexity while preserving accuracy. Kimi k1.5 (Kimi Team et al., 2025) integrates length-based rewards to discourage verbose reasoning. Similarly, O1-Pruner (Luo et al., 2025) detects “length disharmony” and applies harmonizing penalties that promote brevity without sacrificing solution quality. Arora et al. (Arora & Zanette, 2025) use reinforcement learning to train models that dynamically allocate computational resources based on task difficulty, balancing cost and precision. DAST (Shen et al., 2025b) proposes a Token Length Budget metric that aligns task complexity with output length, encouraging efficiency through targeted penalties and rewards. IBPO (Yu et al., 2025) adopts a constrained RL framework to control the distribution of reasoning across response groups based on inference cost. MRT (Qu et al., 2025) applies meta-reinforcement learning to balance exploration of novel reasoning paths with the exploitation of concise, proven ones. Recent approaches also explore interactive or user-directed mechanisms for length control. Claude 3.7 (Anthropic, 2025), the first hybrid reasoning model, introduces an extended thinking mode where users can prescribe token budgets. L1 (Aggarwal & Welleck, 2025) generalizes this idea with Length Controlled Policy Optimization (LCPO), enabling fully configurable CoT lengths at inference time.
3.2 Implicit Latent CoT
Implicit latent CoT methods enhance token efficiency by shifting reasoning from explicit tokens to latent tokens, encoding reasoning in hidden layers rather than natural language.
A line of knowledge distillation methods (Deng et al., 2023; 2024; Shen et al., 2025c) trains student models to infer the teacher’s internal CoT representations rather than mimic explicit token sequences, enabling “vertical” reasoning across transformer layers. Chain of Continuous Thought (COCONUT)(Hao et al., 2024) replaces token-level reasoning chains with autoregressively generated latent embeddings, which are fed back into the model to emulate breadth-first search during problem-solving. Compressed CoT (CCoT)(Cheng & Van Durme, 2024) introduces contemplation tokens—dense, compressed representations of full reasoning chains—significantly reducing inference latency while maintaining accuracy.
Heima(Shen et al., 2025a) condenses CoT stages into latent thinking tokens and incorporates an explanatory prompt at the decoder stage to interpret the compressed reasoning. SoftCoT (Xu et al., 2025c) utilizes a small instruction-tuned 1B model to obtain instance-specific latent thought tokens and trains a projection layer to incorporate thought tokens into LLM input. Token-Assorted CoT(Su et al., 2025) mixes latent and text tokens, encoding the initial part of the CoT into VAE-based discrete latent tokens while preserving the remainder as natural language, resulting in a hybrid representation that enhances reasoning efficiency.
While their implementations vary, these approaches share a common goal: optimizing inference by internalizing the reasoning process. Empirical results suggest that implicit latent CoT models can match or even surpass explicit CoT methods in reasoning accuracy while significantly reducing generation costs, demonstrating their scalability and efficiency.
Types Methods Setting Accurracy Model Token Cost
Explicit Compact CoT CoD (Xu et al., 2025b) zero-shot 84.40% GPT-4o 76.40
zero-shot 65.50% Claude 3.5 Sonnet 73.70
few-shot 91.10% GPT-4o 43.90
few-shot 91.40% Claude 3.5 Sonnet 39.80
TALE (Han et al., 2024) zero-shot, prompt 84.46% GPT-4o-mini 77.26
zero-shot, SFT 74.11% LLaMA-3.1-8B-Instruct 149.93
zero-shot, DPO 78.41% LLaMA-3.1-8B-Instruct 113.41
C3oT (Kang et al., 2024) zero-shot 36.92% LLaMA-2-Chat-7B -
zero-shot 47.10% LLaMA-2-Chat-13B -
TokenSkip (Xia et al., 2025) zero-shot, ratio=0.5 86.70% LLaMA-3.1-8B-Instruct 113.05
zero-shot, ratio=0.6 86.10% LLaMA-3.1-8B-Instruct 198.01
zero-shot, ratio=0.7 84.30% LLaMA-3.1-8B-Instruct 169.89
zero-shot, ratio=0.8 82.50% LLaMA-3.1-8B-Instruct 150.12
zero-shot, ratio=0.9 81.10% LLaMA-3.1-8B-Instruct 129.38
zero-shot, ratio=1.0 78.20% LLaMA-3.1-8B-Instruct 113.05
LightThinker (Zhang et al., 2025) zero-shot,tho. 90.14% DeepSeek-R1-Distill-Qwen-7B -
zero-shot,token 87.11% DeepSeek-R1-Distill-Qwen-7B -
zero-shot,tho. 88.25% DeepSeek-R1-Distill-LLaMA-8B -
zero-shot,tho. 85.52% DeepSeek-R1-Distill-LLaMA-8B -
SF (Munkhbat et al., 2025) zero-shot 76.72% DeepSeekMath-7B 184.13
O1-Pruner (Luo et al., 2025) few-shot 96.50% QwQ-32B 343.00
Implicit Latent CoT ICoT-KD (Deng et al., 2023) zero-shot 45.00% GPT-2 Medium -
CODI (Shen et al., 2025c) zero-shot 55.60% LLaMA-3.2-1B -
ICoT-SI (Deng et al., 2024) zero-shot 51.00% Mistral 7B -
COCONUT (Hao et al., 2024) zero-shot 34.10% GPT-2 8.20
CCoT (Cheng & Van Durme, 2024) zero-shot 31.50% LLaMA2-7B-Chat -
Token assorted (Su et al., 2025) zero-shot 37.20% LLaMA-3.1-8B -
SoftCoT (Xu et al., 2025c) zero-shot 85.81% Qwen2.5-7B-Instruct -
Table 2: Benchmarking on recent reasoning efficient methods on GSM8K dataset.
4 Empirical Analyses
This section conducts empirical analyses of the existing reasoning efficient methods from the perspectives of performance and token efficiency. We summarize the used benchmarks of the existing methods and categorize them into 10 reasoning scenarios as follows. (1) Mathematical Reasoning includes GSM8K (Cobbe et al., 2021), GSM8K-Zero (Chiang & Lee, 2024), SVAMP (Patel et al., 2021), AQuA Ling et al. (2017), ASDiv (Miao et al., 2020), MathBench (Liu et al., 2024b), TheoremQA (Chen et al., 2023), MATH (Hendrycks et al., 2021), MathQA Yu et al. (2023), AIME24 (Olympiad, 2025), Olympiad-Bench (He et al., 2024a), GPQA (Rein et al., 2024). (2) Causal Reasoning includes QASC (Khot et al., 2019), WorldTree (Jansen et al., 2018) (3) Code Reasoning includes LiveCodeBench (Jain et al., 2024), Codeforces, SWE-bench (Jimenez et al., 2023). (4) Logical Reasoning includes ProntoQA (Saparov & He, 2023), LogiQA (Liu et al., 2020), Reclor (Yu et al., 2020). (5) Symbolic Reasoning includes CoinFlip (Wei et al., 2022). (6) Commonsense Reasoning includes CommonsenseQA (Talmor et al., 2019), OpenbookQA (Mihaylov et al., 2018), ECQA (Aggarwal et al., 2021), StrategyQA (Geva et al., 2021). (7) General Reasoning includes BIG-Bench (Srivastava et al., 2022), BIG-Bench Hard (Suzgun et al., 2022), HotPotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), MMLU (Hendrycks et al., 2020), MMMLU (Yue et al., 2024), ScienceQA (Lu et al., 2022), SciBench (Wang et al., 2024c). (8) Visual Reasoning MMMU (Yue et al., 2024), MATH-Vision (Wang et al., 2024a), MathVista (Lu et al., 2024) (9) Agent Reasoning includes TAU-bench (Yao et al., 2024), Keys-Finding Maze (Su et al., 2025). (10) Task-specific Reasoning includes PubMedQA (Jin et al., 2019). Then, we demonstrate the performance and token costs of the existing methods on one common dataset GSM8K (Cobbe et al., 2021). We list these results in Table 2.
5 Limitations and Challenges
We discuss the limitations and challenges of the existing reasoning efficient methods from the perspectives of user experience, interpretability, safety, and application.
5.1 User-centric Controllable Reasoning
Recent advancements in LRMs, such as OpenAI’s o3 (OpenAI, 2025) and Anthropic’s Claude 3.7 (Anthropic, 2025), have introduced user-configurable reasoning modes, allowing users to choose whether the model engages in explicit reasoning or provides direct answers. Additionally, these models enable users to control the complexity and length of the reasoning process, adapting to different needs and preferences.
This control level is particularly useful in diverse applications, e.g., in educational settings, users may prefer step-by-step explanations for questions, whereas in real-time decision-making tasks, concise responses are more desirable. Allowing users to adjust reasoning depth enables LRMs to balance efficiency and transparency, enhancing user experience.
Future research should explore more refined control mechanisms, such as interactive reasoning settings that dynamically adjust based on user feedback. Besides, developing personalized reasoning profiles could allow LRMs to learn and adapt to individual preferences over time, providing a balance between reasoning depth, speed, and interpretability.
5.2 Trade-off Between Interpretability and Efficiency of Reasoning
Compared to LLMs, LRMs offer significantly better interpretability due to their structured reasoning process. By explicitly generating intermediate reasoning steps, LRMs allow users to trace how a conclusion is reached, making them particularly valuable for applications where transparency and verifiability are critical, such as scientific research (Rane et al., 2023), medical diagnosis (Ullah et al., 2024), and legal decision-making (Cheong et al., 2024). However, current efficiency-focused LRMs may compromise this interpretability. Many recent methods designed to accelerate LRM inference reduce the number of explicit reasoning steps or shift reasoning to latent representations, making it harder to understand how a model arrives at its conclusions.
Also, the importance of interpretability varies depending on the application. In domains such as healthcare and legal reasoning, where explanations are essential for accountability and human oversight, explicit reasoning steps are preferred despite their computational cost. Conversely, in real-time decision-making tasks, such as automated trading or robotics, efficiency often takes precedence over transparency, making implicit reasoning more desirable. Hybrid approaches, which dynamically adjust the level of explicit reasoning based on task complexity, offer a potential solution but require further refinement to prevent critical reasoning steps from being lost in the pursuit of efficiency.
To address this trade-off more effectively, future research should focus on developing adaptive inference strategies that optimize the balance between reasoning efficiency and interpretability. One promising direction is the integration of external verification mechanisms, such as symbolic reasoning (Besold et al., 2021; Gaur & Saunshi, 2023; Sui et al., 2024) or retrieval-based justifications (Gao et al., 2023), which can provide post-hoc explanations for implicit reasoning models. Besides, new empirical studies are needed to quantify how different efficiency techniques impact both model accuracy and human trust, guiding the development of LRMs that are both efficient and interpretable in real-world scenarios.
5.3 Ensuring Safety of Efficient Reasoning
Although the existing methods improve the token efficiency of the LRMs, they may destroy the alignment of LRMs, increasing the potential safety risks, e.g., jailbreaking attacks (Liu et al., 2024d; He et al., 2025a) and privacy leakage (Li et al., 2023).
Firstly, the current training-based token-efficient methods either train the LRMs to prefer shorter generations (Kang et al., 2024; Han et al., 2024) or adopt RL and incentivize concise responses via rule-based reward (Qu et al., 2025; Luo et al., 2025; Kimi Team et al., 2025). Since the safety alignment is conducted on the original long reasoning generations and the safety of the shorter reasoning generations can not be guaranteed, these training processes might break the safety alignment of the original LRMs.
Secondly, as one piece of evidence, researchers (OpenAI, 2025c) found that the frontier LRMs tend to exploit the loopholes once they get a chance. In addition, although they tried to use another LLM to monitor the intermediate CoT, penalizing their misbehavior can not effectively alleviate this problem but further guide them to hide their misconduct intent. From this phenomenon, we suspect that the existing token-efficient methods unintentionally guide the LRMs to hide their harmful intent during the process of making their response more concise, increasing the difficulty of safeguarding LRMs.
To address this problem, one promising direction is to add safety constraints during the training process, like data filtering for the SFT/DPO data and designing the safety-related reward in RL training. Besides, the failure of current monitors may be due to LRMs’ ability being stronger than LLM-based guard models. Thus, it is worth designing stronger reasoning-based safeguard models (Liu et al., 2025b) to monitor the training data or LRMs.
5.4 Broader Application of Efficient Reasoning
As shown in Table 1, existing LRMs are primarily applied in math (Xia et al., 2025; Li et al., 2025), code (Kimi Team et al., 2025), or AI research (OpenAI, 2025a) scenarios.
The first reason is that these tasks have relatively fixed answers, making it easier to construct objectives, e.g., preparing reasoning data, formulating preference loss functions, or rule-based rewards. In contrast, other domains, like social sciences (Manning et al., 2024; Thapa et al., 2025), emotional intelligence (Wu et al., 2025b), creative writing (OpenAI, 2025b), typically involve open-ended questions, making it difficult to formulate clear objectives.
The second reason is that these scenarios, like math or research, are not highly time-sensitive, allowing for more computational resources to be allocated for reasoning and optimization. The high computational demand and latency of LRMs constrain their applicability in broader time-sensitive domains, such as robotic manipulations (Ji et al., 2025; Google, 2025; Nvidia, 2025), financial trading (Ding et al., 2024), autonomous driving (Yang et al., 2023).
However, recently developed efficient reasoning methods (Cheng & Van Durme, 2024; Anthropic, 2025; Kimi Team et al., 2025) help LRMs reduce thinking tokens, optimize timing and memory usage, and thus enhance feasibility in real-time applications. For the open-ended questions, efficient reasoning methods enable LRMs to generate more structured and consistent responses while balancing interpretability and computational cost.
6 How can the inference efficiency of LRMs be further improved?
While the existing methods are effective, we present alternative strategies that could further improve inference efficiency while maintaining the reasoning quality, including new architectures, model merge, and agent routers.
6.1 New Architecture
Hybrid Autoregressive and Diffusion Models. The fundamental limitation of autoregressive models is their sequential nature, which makes inference slow, particularly for reasoning tasks that require long chains of intermediate steps. A potential solution is integrating diffusion models into LRMs (Nie et al., 2025). Diffusion models generate entire sequences in parallel, allowing for global reasoning structure optimization rather than token-by-token generation. But the challenge lies in controlling the generated reasoning steps to ensure logical consistency. A promising direction is hybrid architectures that use autoregression for fine-grained control over reasoning while leveraging diffusion-based sampling for efficiency, enabling LRMs to reason in a structured yet accelerated manner.
Memory-Efficient Transformer Variants. One of the primary inefficiencies in LRMs stems from the quadratic complexity of self-attention. Applying linear attention mechanisms(e.g., RWKV (Peng et al., 2023)) or state-space models (e.g., Mamba (Gu & Dao, 2023)) could drastically reduce memory consumption and improve inference speed. The challenge is that such architectures often struggle with long-range dependencies, which are crucial for reasoning. A key question is whether hybrid models can selectively apply full attention for critical reasoning steps while using approximate attention elsewhere to optimize efficiency.
Graph-Based Reasoning Models. Autoregressive LRMs process information sequentially, generating one token at a time. While effective, this approach struggles with complex multi-step reasoning tasks where different pieces of information must be retrieved, combined, and reasoned over in a structured manner. Graph-based reasoning models (Besta et al., 2024) offer a structured alternative to autoregressive LRMs by representing reasoning as a graph, where nodes encode intermediate steps and edges define logical dependencies. This enables parallel exploration of multiple reasoning paths, reducing inefficiencies inherent in sequential token generation. Monte Carlo Tree Search (MCTS) (Silver et al., 2017) enhances this by adaptively expanding promising reasoning branches while pruning redundant ones, optimizing both token efficiency and inference time. However, key challenges include designing training objectives for graph construction, optimizing traversal heuristics, and balancing interpretability with efficiency. Future research should explore hybrid architectures that fuse explicit graph reasoning with neural representations (He et al., 2024b; 2025b; Zhou et al., 2023), ensuring both scalability and robust logical coherence.
6.2 Model Merge
The underlying principle of the existing token-efficient methods can be summarized as integrating the strength of the conventional LLMs, i.e., fast responses and low costs, and the strength of the LRMs, i.e., deliberative reasoning and accurate responses.
The existing training-based methods (Han et al., 2024; Luo et al., 2025; Qu et al., 2025) typically involve reasoning data curation and post-training techniques such as SFT, DPO, or RL, making the process complex and expensive. On the other hand, the existing training-free methods (Sui et al., 2025) typically just use promoting engineering to guide the LRMs to save the tokens, limiting the adaptability and effectiveness across diverse reasoning tasks.
To solve this problem, another training-free method model merge (Yang et al., 2024; Wu et al., 2025a) becomes a promising technique. Concretely, we can simply merge the model weights of one conventional LLM and the corresponding LRM to take their advantages together (Kimi Team et al., 2025). During this process, we provide several key points that need to be solved in the future. First, we need to determine which modules or neurons in models should be merged. Should we merge the neurons in shallow networks or deep networks? Then, we should assign merging weights for the merging units. Should we assign static or dynamic weights for each unit? Third, we should consider how to merge models with different architectures and model sizes, e.g., LLaMA-3.1 Instruct 8B (Grattafiori et al., 2024) and DeepSeekR1-Distill-Qwen-7B (Guo et al., 2025).
6.3 Agent Router
Agent routing could further improve efficiency by directing different parts of a query to specialized agents. By routing the query to the most appropriate agent based on task complexity, this strategy would optimize resource usage and enable faster inference, particularly for tasks that require domain-specific knowledge or specialized reasoning.
There are two existing routing strategies: one based on router models and the other on metrics like confidence. Routellm (Ong et al., 2025) introduces several efficient router models that dynamically select either a stronger or a weaker LLM during inference to balance cost and response quality. Self-REF (Chuang et al., 2025b) routes LLMs by training them to reliably express their confidence in the correctness of their answers. Additionally, Chuang et al. (2025a) explores an uncertainty-based small language model (SLM) routing approach, where high-stakes queries are offloaded to more robust LLMs whenever the SLM produces low-confidence responses. These methods offer exciting possibilities for enhancing inference efficiency in LRMs, particularly in reducing computation time and resource use while maintaining high performance across a range of tasks.
7 Final Remarks
This survey provides an overview of efficient inference techniques for large reasoning models, highlighting the challenges and recent advancements in this area. As reasoning models continue to grow in scale, the computational cost of inference becomes a major bottleneck, necessitating methods that improve efficiency while maintaining performance. We categorized existing approaches, discussing their trade-offs and practical implications. We hope this survey provides a foundation for further research in this area, encouraging the development of more effective and computationally feasible reasoning models.